# Teleportation Fidelity of Binary Tree Quantum Repeater Networks

### v2

@@ III.2 Undirected Binary Trees
$l_{d}^{(i)}=\begin{cases}3\cdot 2^{d+\frac{i}{2}-1}-(2d-i+5)\cdot 2^{i-2},&\text{if}\leavevmode\nobreak\ d<i\leq 2d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ i\leavevmode\nobreak\ \text{is even}\\ 2^{d+\frac{i+1}{2}}-(2d-i+5)\cdot 2^{i-2},&\text{if}\leavevmode\nobreak\ d<i\leq 2d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ i\leavevmode\nobreak\ \text{is odd}\\ 3\cdot 2^{d+\frac{i}{2}-1}-(i+3)\cdot 2^{i-2},&\text{if}\leavevmode\nobreak\ i\leq d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ i\leavevmode\nobreak\ \text{is even}\\ 2^{d+\frac{i+1}{2}}-(i+3)\cdot 2^{i-2},&\text{if}\leavevmode\nobreak\ i\leq d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ i\leavevmode\nobreak\ \text{is odd}\\ 0,&\text{otherwise}.\end{cases}$
Like the previous cases, here we can also express Eq. (15) in terms of NN where $N=2^{d+1}-1$. The full derivation of the formula is given in appendix A.2.
IV Quantum Advantage in Binary Tree Networks
@@
There are various kinds of comprehension for the concept of quantum advantage in the distributed teleportation of a repeater network. The $F^{\text{tel}}_{\text{avg}}$ has been plotted with the input state parameter in Figs. 3(a) and 3(b) to determine the value of $p$ at which each of these four networks exhibits quantum advantages. The quantum advantage is demonstrated by taking into account all types of trees that have the same number of nodes, meaning that depth $d$ differs for symmetric and asymmetric trees. We presented the results for $N=15$ in Fig. 3(a) and for $N=127$ in Fig. 3(b). For all values of $p$ beyond $0.638$, $0.511$, $0.699$, and $0.698$, respectively, DABT, DSBT, UABT, and USBT exhibit quantum advantage shown in Fig. 3(a). The values are indicated in the diagram by double circle markers (see dashed horizontal line). The quantum advantage is also obtained for a range of $p$ larger than $0.931$, $0.666$, $0.936$, and $0.868$, respectively, for $N=127$ (Fig. 3(b), dashed horizontal line).
For $N=15$, given that directed binary trees include fewer pathways than undirected binary trees, it is intriguing to note that they consistently outperform the latter in terms of permitting a larger range of $p$. The average pathlength $\langle L\rangle$ has values of $3$, $1.83$, $3.45$, and $3.5$ for DABT, DSBT, UABT, and USBT, respectively, for $N=15$. However, for DABT, DSBT, UABT, and USBT, the corresponding values of $\langle L\rangle$ are $21.67$, $3.19$, $22.31$, and $8.35$ in the case of $N=127$. In both circumstances, the DSBT value of $\langle L\rangle$ is at its minimum. It can be shown for large $N$, $\langle L\rangle$, is minimum, compared to other three networks. These observations are consistent with the $F^{\text{tel}}_{\text{avg}}$ values of all the networks.
@@
Maximally Entangled links : Next, we look into a scenario where we consider some link to be a maximally entangled state. In reference [1], the role of maximally entangled states in determining the average fidelities of loopless topologies like start, chain and branching trees. The assignment of maximally entangled states is also going to play a crucial role in determining the average fidelity. In Fig. 5(a), 5(b), 5(c), and 5(d), we have used contour plot to show how $F^{\text{tel}}_{\text{avg}}$ changes as a function of the number of maximally entangled states ($M$) vs $p$. The pink region in the figures indicate the region where $F^{\text{tel}}_{\text{avg}}<\frac{2}{3}$. The cyan region in the figures indicates the ranges of $M$ and $p$ for which the network teleportation fidelity is $F^{\text{tel}}_{\text{avg}}>\frac{2}{3}$. It is clear from each graph that for a lower value of $p$, we require a larger number of maximally entangled states in each of these graphs for $F^{\text{tel}}_{\text{avg}}$ to cross the threshold limit $2/3$. As expected for a larger value of $p$, we do not require maximally entangled states for $F^{\text{tel}}_{\text{avg}}$ to cross the threshold limit. For $N=15$, the threshold value of $M$ required for quantum advantage are $4$, 77, 88, and $8$ for DSBT, DABT, UABT, and USBT. Among the four graphs, DSBT performs the best as there is a larger range of $p$ for which there is a quantum advantage in the network.
@@
Figure 5: The contour plot of $F^{\text{tel}}_{\text{avg}}$: (a) DABT, (b) UABT. In case of $p=0.333$, the network achieves $F^{\text{tel}}_{\text{avg}}=2/3$ when $M=7$ and 8 for DABT and UABT, respectively. The contour plot of $F^{\text{tel}}_{\text{avg}}$: (c) DSBT, and (d) USBT. In case of $p=0.333$, the network achieves $F^{\text{tel}}_{\text{avg}}=2/3$ when $M=4$ and 8 for DSBT and USBT, respectively.
@@
Large $N$ limit: For DABT, we have analytically and numerically demonstrated $F^{\text{tel}}_{\text{avg}}$ reaches $1/2$ for a large $d$ (large $N$). In Eq. (18), and the Fig. 12 of appendix A.2, the additional term $\varepsilon$ is plotted as a function of $N(d)$ that decreases to zero as the number of nodes in the tree increases. This is also true for UABT, USBT, and DSBT (see the Eqns. (29), (25), and Fig. 12 in the appendix). Clearly, regardless of the network configuration, the plot (in the appendix A.2) indicates that additional parts added with $1/2$ converge to zero for a large $N(d)$ limit. In case of DSBT, the additional part $\varepsilon(N)=\mathcal{O}\left(1/\log_{2}N\right)$ (See appendix A.1) and thus it decays slowly with $N(d)$ compared to other binary trees. Hence we get a larger additional term in the case of DSBT. For numerical verification, we computed the $F^{\text{tel}}_{\text{avg}}$ with increasing $N$ for various $p$ values, namely $p=0.333$ and $p=0.9$, in order to see the quantum advantage in the large $N$ limit. Alongside this, we have also plotted for $p=1/2$ and $1/\sqrt{2}$ to verify the formula with the numerical result. The results are reported in Figs.  6(a)($p=0.333$), 6(b)($p=0.9$), 6(c)($p=1/2$), and 6(d)($p=1/\sqrt{2}$). We observe $F^{\text{tel}}_{\text{avg}}$ is converging to $1/2$ but with different rates. It is evident that for low values of $p$ ($p=0.333$), each of these networks $F^{\text{tel}}_{\text{avg}}$ rapidly stabilizes to the value $1/2$. Conversely, for very big NN ($\sim 10^{8}$), DSBT exhibits quantum advantage at $p=0.9$. Nevertheless, inside the $10^{3}$ network size, DABT, UABT, and USBT lose the quantum advantage. Keep in mind that we wish to display comparison outcomes from four networks here. For this, each graph’s depth $d$ is selected so that every network has the same size.
@@
V Average teleportation fidelity for uniform distribution
In a real network (based on photonic qubits), parameters like detector efficiency, the likelihood of a photon being emitted in the correct mode, the probability of a photon being successfully sent into the fiber, attenuation losses in the fiber, etc., cause the qubit’s coherence to be lost [75]. This loss also varies from node to node, as it depends upon the distance between the nodes and also on the individual capacity of repeaters. This variation can be modeled by different Werner state parameters $p_{i}$ for each link [75]. For a fixed number of nodes $N$ in each class of binary trees, we assign weights $\{p_{i}\}$ to the edges, where each weight is independently drawn from a uniform distribution, $p_{i}\sim\mathcal{U}(0,1)$. Since a binary tree with $N$ nodes contains N1N-1 edges, each tree configuration is associated with N1N-1 such weights. To include statistical robustness, we repeat this random weight assignment process 100 times ($T_{1},T_{2},...,T_{100}$) for each tree instance, and each time, the corresponding average teleportation fidelity, $F^{\text{tel}}$. We then take the average of these values across all trials to obtain $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$. The result is shown as a histogram in Fig. 7.
@@ V Average teleportation fidelity for uniform distribution
Nodes DABT DSBT UABT USBT
$F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$ $F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$ $F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$ $F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$
@@
VI Conclusion
This article presents a comprehensive analytical investigation into the ability of hierarchical binary networks to transmit quantum information via teleportation. We consider four distinct network topologies—the directed and undirected asymmetric binary trees (DABT, UABT) and the directed and undirected symmetric binary trees (DSBT, USBT)—to understand the influence of directionality and symmetry on network performance. The quality of quantum information transfer is quantified by the average of the maximum teleportation fidelity, $F^{\text{tel}}_{\text{avg}}$. By modeling the network edges with simple Werner states, we derive analytical expressions for $F^{\text{tel}}_{\text{avg}}$ for each network type and explore their limiting values as the number of nodes, $N$, approaches infinity. This analytical framework allows us to precisely identify the parameter ranges for which these networks can exhibit a quantum advantage, defined by $F^{\text{tel}}_{\text{avg}}>2/3$. Furthermore, we explore the strategic role of adding maximally entangled states to the network to enhance this quantum advantage. To bridge the gap between idealized models and real-world systems, we also investigate the average teleportation fidelity in a more general setting where link parameters are heterogeneous and chosen from a uniform distribution. Our work provides a rigorous foundation for identifying resourceful tree networks for quantum information transmission (teleportation). The analytical expressions obtained, along with the detailed exploration of large-$N$ convergence and the roles of symmetry and directionality, contribute to a deeper understanding of quantum network topologies. In this process, we provide a generalized method for the analytical calculation of pathlengths in all four graph types, a methodological advancement that can be applied to more general Cayley tree structures.
Acknowledgment : S.R. would like to acknowledge DST-India for the INSPIRE Fellowship (IF220695) support. Authors also acknowledge Kishore Kothapalli for having useful discussions, which helps to improve the quality of the manuscript.
@@ A.1 Directed Binary Trees
$l_{d}^{(r)}=\begin{cases}2^{d+1}-2^{r},&\text{if}\leavevmode\nobreak\ r\leq d\\ 0,&\text{otherwise}.\end{cases}$ (23)
@@
$=\begin{cases}\displaystyle\frac{1}{2}+\frac{2^{d+1}-d-2}{2\left[(d-1)\cdot 2^{d+1}+2\right]},&\text{if }\leavevmode\nobreak\ p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p\left[(2^{d}-1)-p(2^{d+1}-1)+2^{d}p^{d+1}\right]}{2(1-p)(1-2p)\left[(d-1)\cdot 2^{d}+1\right]},&\text{otherwise},\end{cases}$ (24)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{DSBT}}$ $=$ $\begin{cases}\displaystyle\frac{1}{2}+\frac{N-\log_{2}(N+1)}{2\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{if }\leavevmode\nobreak\ p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p[N(1-2p)+(N+1)p^{\log_{2}(N+1)}-1]}{4(1-p)(1-2p)\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{otherwise}.\end{cases}$ (26)

### v3

@@ 3.3 Undirected Binary Trees
$l_{d}^{(i)}=\begin{cases}3\cdot 2^{d+\frac{i}{2}-1}-(2d-i+5)\cdot 2^{i-2},&\text{if}~d<i\leq 2d~\text{and}~i~\text{is even}\\ 2^{d+\frac{i+1}{2}}-(2d-i+5)\cdot 2^{i-2},&\text{if}~d<i\leq 2d~\text{and}~i~\text{is odd}\\ 3\cdot 2^{d+\frac{i}{2}-1}-(i+3)\cdot 2^{i-2},&\text{if}~i\leq d~\text{and}~i~\text{is even}\\ 2^{d+\frac{i+1}{2}}-(i+3)\cdot 2^{i-2},&\text{if}~i\leq d~\text{and}~i~\text{is odd}\\ 0,&\text{otherwise}.\end{cases}$
Like the previous cases, here we can also express Eq. (14) in terms of NN where $N=2^{d+1}-1$. The full derivation of the formula is given in appendix A.2. Similarly, for the both undirected binary trees, $F^{\text{tel}}_{\text{avg}}(N,p)=1$ when $p=1$. In the case of FT, the same formula holds.
4 Quantum Advantage in Binary Tree Networks
@@
There are various kinds of comprehension for the concept of quantum advantage in the distributed teleportation ( i.e. teleportation between all possible sources and targets) of a repeater network. The $F^{\text{tel}}_{\text{avg}}$ has been plotted with the input state parameter in Figs. 5(a) and 5(b) to determine the value of $p$ at which each of these four networks exhibits quantum advantages. The quantum advantage is demonstrated by taking into account all types of trees that have the same number of nodes, meaning that depth $d$ differs for symmetric and asymmetric trees. We presented the results for $N=15$ in Fig. 5(a) and for $N=127$ in Fig. 5(b). For all values of $p$ beyond $0.638$, $0.511$, $0.699$, and $0.698$, respectively, DABT, DSBT, UABT, and USBT exhibit quantum advantage shown in Fig. 5(a). The values are indicated in the diagram by double-circle markers (see the dashed horizontal line). The quantum advantage is also obtained for a range of $p$ larger than $0.931$, $0.666$, $0.936$, and $0.868$, respectively, for $N=127$ (Fig. 5(b), dashed horizontal line).
For $N=15$, given that directed binary trees include fewer pathways than undirected binary trees, it is intriguing to note that they consistently outperform the latter in terms of permitting a larger range of $p$. The average pathlength $\langle L\rangle$ has values of $3$, $1.83$, $3.45$, and $3.5$ for DABT, DSBT, UABT, and USBT, respectively, for $N=15$. However, for DABT, DSBT, UABT, and USBT, the corresponding values of $\langle L\rangle$ are $21.67$, $3.19$, $22.31$, and $8.35$ in the case of $N=127$. In both circumstances, the DSBT value of $\langle L\rangle$ is at its minimum. It can be shown that for large $N$, $\langle L\rangle$, is minimum, compared to the other three networks. These observations are consistent with the $F^{\text{tel}}_{\text{avg}}$ values of all the networks.
@@
Maximally Entangled links : Next, we look into a scenario where we consider some link to be a maximally entangled state. In reference [19], the role of maximally entangled states in determining the average fidelities of loopless topologies like start, chain and branching trees. The assignment of maximally entangled states is also going to play a crucial role in determining the average fidelity. In Fig. 7(a), 7(b), 7(c), and 7(d), we have used contour plot to show how $F^{\text{tel}}_{\text{avg}}$ changes as a function of the number of maximally entangled states ($M$) vs $p$. The pink region in the figures indicate the region where $F^{\text{tel}}_{\text{avg}}<\frac{2}{3}$. The cyan region in the figures indicates the ranges of $M$ and $p$ for which the network teleportation fidelity is $F^{\text{tel}}_{\text{avg}}>\frac{2}{3}$. It is clear from each graph that for a lower value of $p$, we require a larger number of maximally entangled states in each of these graphs for $F^{\text{tel}}_{\text{avg}}$ to cross the threshold limit $2/3$. As expected for a larger value of $p$, we do not require maximally entangled states for $F^{\text{tel}}_{\text{avg}}$ to cross the threshold limit. For $N=15$, the threshold value of $M$ required for quantum advantage are $4$, 77, 88, and $8$ for DSBT, DABT, UABT, and USBT. Among the four graphs, DSBT performs the best as there is a larger range of $p$ for which there is a quantum advantage in the network.
@@
Figure 7: The contour plot of $F^{\text{tel}}_{\text{avg}}$: (a) DABT, (b) UABT, (c) DSBT, and (d) USBT. In the case of $p=0.333$, the network achieves $F^{\text{tel}}_{\text{avg}}=2/3$ when $M=7$ and 8, 4 and 8 for DABT, UABT, DSBT, USBT, respectively.
@@
Large $N$ limit: For DABT, we have analytically and numerically demonstrated $F^{\text{tel}}_{\text{avg}}$ reaches $1/2$ for a large $d$ (large $N$). In Eq. (18), and the Fig. 14 of appendix A.2, the additional term $\varepsilon$ is plotted as a function of $N(d)$ that decreases to zero as the number of nodes in the tree increases. This is also true for UABT, USBT, and DSBT (see the Eqns. (29), (25), and Fig. 14 in the appendix). Clearly, regardless of the network configuration, the plot (in the appendix A.2) indicates that additional parts added with $1/2$ converge to zero for a large $N(d)$ limit. In case of DSBT, the additional part $\varepsilon(N)=\mathcal{O}\left(1/\log_{2}N\right)$ (See appendix A.1) and thus it decays slowly with $N(d)$ compared to other binary trees. Hence, we get a larger additional term in the case of DSBT. For numerical verification, we computed the $F^{\text{tel}}_{\text{avg}}$ with increasing $N$ for various $p$ values, namely $p=0.333$ and $p=0.9$, in order to see the quantum advantage in the large $N$ limit. Alongside this, we have also plotted for $p=1/2$ and $1/\sqrt{2}$ to verify the formula with the numerical result. The results are reported in Figs. 8(a)($p=0.333$), 8(b)($p=0.9$), 8(c)($p=1/2$), and 8(d)($p=1/\sqrt{2}$). We observe $F^{\text{tel}}_{\text{avg}}$ is converging to $1/2$ but with different rates. It is evident that for low values of $p$ ($p=0.333$), each of these networks $F^{\text{tel}}_{\text{avg}}$ rapidly stabilizes to the value $1/2$. Conversely, for very big NN ($\sim 10^{8}$), DSBT exhibits quantum advantage at $p=0.9$. Nevertheless, inside the $10^{3}$ network size, DABT, UABT, and USBT lose the quantum advantage. Note that we aim to display comparison results from four networks here. For this, the depth $d$ of each graph is selected so that every network has the same size. Moreover, in case of large $N$, the USBT turns into a FT and hence $F^{\text{tel}}_{\text{avg}}$ converges to the value $1/2$ in case of a FT. As depicted, it will not retain quantum advantage up to very large $N$ compared to DSBT.
@@
5 Average teleportation fidelity for uniform distribution
In a real network (based on photonic qubits), parameters like detector efficiency, the likelihood of a photon being emitted in the correct mode, the probability of a photon being successfully sent into the fiber, attenuation losses in the fiber, etc., cause the qubit’s coherence to be lost. This loss also varies from node to node, as it depends upon the distance between the nodes and also on the individual capacity of repeaters. This variation can be modeled by different Werner state parameters $p_{i}$ for each link [83]. Accordingly, the probability that a single transmission attempt succeeds can be written as
$p_{\mathrm{noise}}=p_{\mathrm{det}}\times 10^{-\alpha d/10},$ (15)
where $\alpha$ is the fibre attenuation coefficient and $d$ is the transmission distance. The time required to have one attempt to transfer a qubit is $\tau=d/c$ with $c$ is the speed of light in the fibre. If the photon is successfully detected, then $p_{\mathrm{det}}=1$. In this scenario, $F^{\text{tel}}_{\text{avg}}$ goes down if we increase the intra-node distance ($d$) [19]. For a fixed number of nodes $N$ in each class of binary trees, we assign weights $\{p_{i}\}$ to the edges, where each weight is independently drawn from a uniform distribution, $p_{i}\sim\mathcal{U}(0,1)$. Since a binary tree with $N$ nodes contains N1N-1 edges, each tree configuration is associated with N1N-1 such weights. To include statistical robustness, we repeat this random weight assignment process 100 times ($T_{1},T_{2},...,T_{100}$) for each tree instance, and each time, the corresponding average teleportation fidelity, $F^{\text{tel}}$. We then take the average of these values across all trials to obtain $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$. The result is shown as a histogram in Fig. 9.
@@ 5 Average teleportation fidelity for uniform distribution
DABT DSBT UABT USBT
Nodes $F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$ $F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$ $F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$ $F_{\text{avg}}^{\text{tel}}$ $\langle F_{\text{avg}}^{\text{tel}}\rangle_{T_{1},...,T_{100}}$
@@
6 Conclusion
This article presents a comprehensive analytical investigation into the ability of hierarchical binary networks to transmit quantum information via teleportation. We consider four distinct network topologies - the directed and undirected asymmetric binary trees (DABT, UABT) and the directed and undirected symmetric binary trees (DSBT, USBT) - to understand the influence of directionality and symmetry on network performance. The quality of quantum information transfer is quantified by the average of the maximum teleportation fidelity, $F^{\text{tel}}_{\text{avg}}$. By modeling the network edges with simple Werner states, we derive analytical expressions for $F^{\text{tel}}_{\text{avg}}$ for each network type and explore their limiting values as the number of nodes, $N$, approaches infinity (similar to the fractal tree). This analytical framework allows us to precisely identify the parameter ranges for which these networks can exhibit a quantum advantage, defined by $F^{\text{tel}}_{\text{avg}}>2/3$. Furthermore, we explore the strategic role of adding maximally entangled states to the network to enhance this quantum advantage. To bridge the gap between idealized models and real-world systems, we also investigate the average teleportation fidelity in a more general setting where link parameters are heterogeneous and chosen from a uniform distribution. There are numerous techniques to reduce the loss and decoherence in the fibres, such as using low-loss fibre, improving coupling and optics, and stabilising the phase and polarization. Alongside this, the quantum error correction techniques, such as the Quantum Parity Code (QPC) [84, 85], and the Calderbank-Shor-Steane (CSS) code [86], can be used to improve the communication efficiency due to the resilience to photon loss. Our work provides a rigorous foundation for identifying resourceful tree networks for quantum information transmission (teleportation). The analytical expressions obtained, along with the detailed exploration of large-$N$ convergence and the roles of symmetry and directionality, contribute to a deeper understanding of quantum network topologies. In this process, we provide a generalized method for the analytical calculation of pathlengths in all four graph types, a methodological advancement that can be applied to more general Cayley tree structures.
[…]
Acknowledgement S.R. would like to acknowledge DST-India for the INSPIRE Fellowship (IF220695) support. The authors also acknowledge Kishore Kothapalli for having useful discussions, which helped to improve the quality of the manuscript.
@@ A.1 Directed Binary Trees
$l_{d}^{(r)}=\begin{cases}2^{d+1}-2^{r},&\text{if}~r\leq d\\ 0,&\text{otherwise}.\end{cases}$ (23)
@@
$=\begin{cases}\displaystyle\frac{1}{2}+\frac{2^{d+1}-d-2}{2\left[(d-1)\cdot 2^{d+1}+2\right]},&\text{if }~p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p\left[(2^{d}-1)-p(2^{d+1}-1)+2^{d}p^{d+1}\right]}{2(1-p)(1-2p)\left[(d-1)\cdot 2^{d}+1\right]},&\text{otherwise},\end{cases}$ (24)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{DSBT}}$ $=$ $\begin{cases}\displaystyle\frac{1}{2}+\frac{N-\log_{2}(N+1)}{2\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{if }~p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p[N(1-2p)+(N+1)p^{\log_{2}(N+1)}-1]}{4(1-p)(1-2p)\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{otherwise}.\end{cases}$ (26)